OpenStudy (anonymous):

HELP ME PLEASE Solve for x: −3(x + 3) = −3(x + 1) − 5.

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):

Let's solve your equation step-by-step. −3(x+3)=−3(x+1)−5 Step 1: Simplify both sides of the equation. −3(x+3)=−3(x+1)−5 Simplify: Let's simplify step-by-step. −3(x+3)=−3(x+1)−5 (−3)(x)+(−3)(3)=(−3)(x)+(−3)(1)+−5(Distribute) −3x+−9=−3x+−3+−5 −3x−9=(−3x)+(−3+−5)(Combine Like Terms) −3x−9=−3x+−8 −3x−9=−3x−8 Step 2: Add 3x to both sides. −3x−9+3x=−3x−8+3x −9=−8 Step 3: Add 9 to both sides. −9+9=−8+9 0=1

OpenStudy (texaschic101):

when you get an answer such as 0 = 1, then there is no solution

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you! @ToriaJ i kind of get the idea of how to solve it now

OpenStudy (anonymous):

Okay. Good! :)